# Recalibrating binary probabilistic classifiers

### v1

@@ 4.5 Quasi moment matching based on parametrised receiver operating characteristics
There is no guarantee that the QMM approach presented in Section 4.4 is feasible in the sense that there exists a solution to equation system (7c) or that the solution is unique. This reservation motivates the following approach where instead of beginning with representation (7a) of the posterior probabilities, the starting point is a parametrised representation of the receiver operating characteristic (ROC) curve associated with the target distribution $Q(X,Y)$.
@@
For $AUC_{S}$ as defined in Appendix 0.A.1 below, (8a) implies $AUC_{S}=\Phi\bigl{(}\frac{c}{\sqrt{2}}\bigr{)}$. Hence, if $AUC_{S}$ is known the parameter $c$ of (8b) is determined by
@@
The score $S$ in (8b) can be chosen as $\eta_{P}(X)$ or any probabilistic classifier which approximates $\eta_{P}(X)$. The target prior probability $q$ is known by general assumption for this paper. Similarly to Section 4.4, for QMM to work one has to make the assumption that a prudent choice of $AUC_{S}$ under the target distribution (defined as implied AUC by (10b) and (10a) below) is informed by $AUC_{S}$ observed in the training dataset such that by (8c) also parameter $c$ is known.
However, the distribution function $F_{0}$ of the score $S$ conditional on $Y=0$ appearing in (8b) is assumed not to be known under the target distribution since by general assumption for this paper, only the features but not the labels can be observed in the test dataset. Making use of (8b) therefore requires an iterative approach where in each step a refined estimate of the negative class-conditional score distribution $F_{0}$ is calculated. The numerical example of Section 5 below suggests that such an iterative approaches converges as long as the prior probability $q$ is small.
A further issue when making (8b) operational may occur when the class-conditional distribution function $F_{0}$ or any function approximating it takes the value $1$. In particular, this will happen if $F_{0}$ is approximated by an empirical distribution function. Then the term $\Phi^{-1}\bigl{(}F_{0}(s)\bigr{)}$ is ill-defined. See Appendix 0.A.3 for a workaround to deal with this issue.
Deploying a discrete distribution $F_{0}$ in (8b) as for instance an empirical approximation, makes it unlikely if not impossible to exactly match a pre-defined $AUC_{S}$ when using (8b). To control for the in this case unavoidable deviation from the $AUC_{S}$ objective we take recourse to the original QMM approach as proposed by Tasche [20]. Define
@@ 5 Example
The example presented in this section illustrates the impact of the recalibration methods and assumptions discussed in Section 4 on certain characteristics of the target distribution:
@@
  • The mean of the square root of the class 1 posterior probabilities as an expample of a concave functional of the posterior probabilities.
For the example we chose discrete source and target distributions of a feature with values in an ordered set with 17 elements. These distributions may be interpreted as empirical distributions of samples with many ties or as the genuine distributions of discrete scores or ratings. For instance, the major credit rating agencies Standard & Poor’s, Moody’s and Fitch use rating scales with 17 to 19 different grades.
@@
The source distribution is specified as follows: (1) The conditional feature distribution for class 0 is a binomial distribution with success probability 0.4, the conditional feature distribution for class 1 is a binomial distribution with success probability 0.55. The number of trials for both binomial distributions is 16. (2) The class 1 prior probability is $p=0.01$.
The target distribution is incompletely specified as follows: (1) The unconditional feature distribution is a binomial distribution with number of trials 16 whose success probability is a Vasicek-distributed random variable with mean 0.3 and correlation parameter 0.3. (2) The target class 1 prior probability is $q=0.05$.
Figure 1 shows [^2] Calculations were performed with R (R Core Team [15]). Details of some more involved calculations are described in Appendices 0.A.2 and 0.A.3. The R-scripts can be downloaded from https://www.researchgate.net/profile/Dirk-Tasche. the source and target unconditional feature distributions. They were intentionally chosen to be quite different such that any distribution shift assumed for the recalibration must be significant.
Figure 2 presents the source posterior probabilities and eight different sets of target posterior probabilities that have been calculated with the methods and assumptions discussed in Section 4. The dots for the probabilities have been linearly interpolated to facilitate reading the figure.
Figure 2: Class 1 posterior probabilities for the example in Section 5. The scale of the vertical axis is logarithmic.
@@
As an example for the application of a concave functional to the target posterior probabilities, we considered the function $C(u)=\sqrt{u}$, $u\in[0,1]$, and its expected value $E_{Q}\bigl{[}C(\eta_{Q}(X))\bigr{]}$ under the target distribution. Note that (2) implies $0.05=q\leq E_{Q}\bigl{[}\sqrt{\eta_{Q}(X)}\bigr{]}\leq\sqrt{q}\approx 0.2236$.
It is clear from column ‘mean(sqrt(probs))’ of Table 1 that AUC is a driver of the value of the concave functional. The lower the value of AUC, the higher is the value of the concave functional. Hence it makes sense to have AUC as a second objective to be matched, in addition to requiring that the target class 1 prior probability is reached. But even if both of these conditions are met there can still be variation in the value of the concave functional. This is demonstrated by the ‘Platt scaling’ posterior probabilities for which $E_{Q}\bigl{[}\sqrt{\eta_{Q}(X)}\bigr{]}$ takes a notedly lower value than for the other four QMM posterior probabilities with almost identical values.
@@ 6 Conclusions
Recalibration of binary probabilistic classifiers to a target prior probability is an important task in areas like credit risk management. This paper presents analyses and methods for recalibration from a distribution shift perspective. It has turned out that distribution shift assumptions linked to the performance in terms of area under the curve (AUC) of a probabilistic classifier under the source distribution are useful for the design of meaningful recalibration methods. Two new methods called parametric CSPD (covariate shift with posterior drift) and ROC-based QMM (quasi moment matching) have been proposed and have been tested together with some other methods in an example setting. The outcomes of this testing exercise suggest that the QMM methods discussed in the paper can provide appropriately conservative results in evaluations with concave functionals like for instance risk weights functions for credit risk.
@@ Appendix 0.A Appendix
0.A.1 Area Under the Curve (AUC)
For the readers’ convenience, we present here the population-level (in contrast to sample-based) representations of AUC (Area Under the Curve333‘Curve’ refers to ROC (Receiver Operating Characteristic). See Fawcett [8] and the references therein for the most common definitions of ROC and AUC.) that are used in the main text of the paper.
$AUC_{S}\ =\ P^{\ast}[S_{1}>S_{0}]+\frac{1}{2}\,P^{\ast}[S_{1}=S_{0}],$ (10a)
However, $AUC_{S}$ can also be determined if the distribution $P(S)$ of $S$ and the posterior probabilities $\eta_{P}(S)=P[Y=1\,|\,S]$ are known. Then with $p=E_{P}[\eta_{P}(S)]$ it holds that
$P_{1}[S\in M]=\frac{E_{P}\bigl{[}\eta_{P}(S)\,\mathbf{1}_{M}(S)\bigr{]}}{p}\ \text{and}\ P_{0}[S\in M]=\frac{E_{P}\bigl{[}(1-\eta_{P}(S))\,\mathbf{1}_{M}(S)\bigr{]}}{1-p}$ (10b)
for all measurable sets $M\subset\boldsymbol{\mathcal{S}}$. See footnote 1 for the definition of $\mathbf{1}_{M}$. Plug $P_{0}$ and $P_{1}$ from (10b) in (10a) to compute $AUC_{S}$.
To indicate the way $AUC_{S}$ is computed, in this paper we refer simply to AUC if $AUC_{S}$ is assumed to be computed or inferred by means of (10a) irrespectively of the origin of $P_{0}$ and $P_{1}$. We refer to implied AUC if $AUC_{S}$ is assumed to be computed by means of (10b) in combination with (10a). See Appendix 0.A.2 for a more detailed formula for implied AUC in the case of a discrete-valued score $S$.
0.A.2 The AUC of discrete auto-calibrated probabilitistic classifiers
How to calculate implied AUC for an auto-calibrated [^4] The probabilistic classifier $S$ is auto-calibrated if for all $s$ in the range of $S$ it holds that $\mu[Y=1\,|\,S=s]=s$ (Tsyplakov [24], Section 2.2). probabilistic classifier $S$ under a joint distribution $\mu(S,Y)$ of $S$ and $Y\in\mathcal{Y}=\{0,1\}$? We assume here that the distribution of $S$ under distribution $\mu$ is discrete, i.e. it is is given by pairs of score values $s_{i}$ and their probabilities $\mu[S=s_{i}]=\pi_{i}>0$, $i=1,\ldots,n$. Then it follows from Appendix 0.A.1 that
@@ 0.A.2 The AUC of discrete auto-calibrated probabilitistic classifiers
0.A.3 Adapting discrete distributions for QMM

### v2

@@ 4.5 Quasi moment matching based on parametrised receiver operating characteristics
There is no guarantee that the QMM approach presented in Section 4.4 is feasible in the sense that there exists a solution to equation system (8c) or that the solution is unique. This reservation motivates the following alternative approach where instead of beginning with representation (8a) of the posterior probabilities, the starting point is a parametrised representation of the receiver operating characteristic (ROC) curve associated with the target distribution $Q(X,Y)$.
@@
For $AUC_{S}$ as defined in Section 3.3, (10a) implies $AUC_{S}=\Phi\bigl{(}\frac{c}{\sqrt{2}}\bigr{)}$. Hence, if $AUC_{S}$ is known the parameter $c$ of (10b) is determined by
@@
The score $S$ in (10b) can be chosen as $\eta_{P}(X)$ or any probabilistic classifier which approximates $\eta_{P}(X)$. The target prior probability $q$ is known by general assumption for this paper. Similarly to Section 4.4, for QMM to work one has to make the assumption that a prudent choice of $AUC_{S}$ under the target distribution (defined as implied AUC by (3b) and (3a) above) is informed by $AUC_{S}$ observed in the training dataset such that by (10c) also parameter $c$ is known.
However, the distribution function $F_{0}$ of the score $S$ conditional on $Y=0$ appearing in (10b) is assumed not to be known under the target distribution since by general assumption for this paper, only the features but not the labels can be observed in the test dataset. Making use of (10b) therefore requires an iterative approach where in each step a refined estimate of the negative class-conditional score distribution function $F_{0}$ is calculated.
Denote by $f$ an unconditional density of $S$ under $Q$ and by $f_{0}$ a density of $F_{0}$. Then the posterior probability $Q[Y=1\,|\,S=s]$ can be represented as
[…]
for $s$ in the range $\mathcal{S}$ of $S$. By (10b), (11a) implies
$f_{0}(s)\ =\ \frac{f(s)}{1-q}\,\left(1-\frac{1}{1+\frac{1-q}{q}\,\exp\bigl{(}c^{2}/2-c\,\Phi^{-1}(F_{0}(s))\bigr{)}}\right),\quad s\in\mathcal{S}.$ (11b)
In the case of scores $S$ under a discrete or empirical distribution, as described in Appendices 0.A.1 and 0.A.2, (11b) can be treated as a fixed point equation for the probabilities $Q[S=s\,|\,Y=0]=f_{0}(s)$ and be solved by a straightforward fixed-point iteration with intial values $Q[S=s]$, $s\in\mathcal{S}$. The numerical example of Section 5 below suggests that such an iterative approach converges as long as the prior probability $q$ is small and, hence, $f$ and $f_{0}$ are close to each other.
A further issue when making (10b) operational may occur when the class-conditional distribution function $F_{0}$ or any function approximating it takes the value $1$. In particular, this will happen if $F_{0}$ is approximated by an empirical distribution function. Then the term $\Phi^{-1}\bigl{(}F_{0}(s)\bigr{)}$ is ill-defined. See Appendix 0.A.2 for a workaround to deal with this issue.
Deploying a discrete distribution $F_{0}$ in (10b) as for instance an empirical approximation, makes it unlikely if not impossible to exactly match a pre-defined $AUC_{S}$ when using (10b). To control for the unavoidable deviation from the $AUC_{S}$ objective in this case we take recourse to the original QMM approach as proposed by Tasche [22]. Define
@@ 5 Example
The example presented in this section illustrates the impact of the recalibration methods and assumptions discussed in Section 4 on the following characteristics of the target distribution:
@@
  • The mean of the square root of the class 1 posterior probabilities as an expample involving a concave function of the posterior probabilities.
For the example we chose discrete source and target distributions of a feature (called score in the following) with values in an ordered set with 17 elements. These distributions may be interpreted as empirical distributions of samples with many ties or as the genuine distributions of discrete scores or ratings. For instance, the major credit rating agencies Standard & Poor’s, Moody’s and Fitch use rating scales with 17 to 19 different grades.
The source distribution is specified as follows: (1) The conditional feature distribution for class 0 is a binomial distribution with success probability 0.4, the conditional feature distribution for class 1 is a binomial distribution with success probability 0.55. The number of trials for both binomial distributions is 16 such that the support of the distribution is the set $\{0,1,\ldots,16\}$. (2) The class 1 prior probability is $p=0.01$.
The target distribution is incompletely specified as follows: (1) The unconditional feature distribution is a binomial distribution with number of trials 16 whose success probability is a Vasicek-distributed random variable (Section 2.2 of Meyer [13]) with mean 0.3 and correlation parameter 0.3. (2) The target class 1 prior probability is $q=0.05$.
Figure 1 shows [^3] Calculations were performed with R (R Core Team [17]). Details of some more involved calculations are described in Appendices 0.A.1 and 0.A.2. The R-scripts can be downloaded from https://www.researchgate.net/profile/Dirk-Tasche. the source and target unconditional score distributions. They were intentionally chosen to be quite different such that any distribution shift assumed for the recalibration must be significant.
@@
Figure 2 presents the source posterior probabilities and eight different sets of target posterior probabilities that have been calculated with the methods and assumptions discussed in Section 4:
  • Capped scaling: Section 4.1
  • Label shift: Section 4.2
  • FJS: Factorizable joint shift, Section 4.3
  • Platt scaling: Section 4.4, Eq. (8b)
  • ROC QMM: ROC-based QMM, Section 4.5, Eq. (10b) and Eq. (10c)
  • 2-param QMM: 2-parameter QMM, Section 4.5, Eq. (12a) and Eq. (12b)
  • Logistic CSPD: Section 4.4, Eq. (8a) with $F(x)=\frac{1}{1+\exp(-x)}$
  • Normal CSPD: Section 4.4, Eq. (8a) with $F(x)=\Phi(x)$, the standard normal distribution function
Figure 2: Class 1 posterior probabilities for the example in Section 5. The scale of the vertical axis is logarithmic. The dots for the probabilities have been connected with straight lines for better readability. ‘Source’ refers to the posterior probabilities in the source distribution without recalibration.
@@
As an example for the application of a concave function to the target posterior probabilities, we chose the function $C(u)=\sqrt{u}$, $u\in[0,1]$, and the related expected value $E_{Q}\bigl{[}C(\eta_{Q}(X))\bigr{]}$ under the target distribution. Note that (2) implies $0.05=q\leq E_{Q}\bigl{[}\sqrt{\eta_{Q}(X)}\bigr{]}\leq\sqrt{q}\approx 0.2236$.
It is clear from column ‘mean(sqrt(probs))’ of Table 1 that AUC is a driver of the mean of the concave function of the posterior probabilities. The lower the value of AUC, the higher is the mean of the concave function of the probabilities. Hence it makes sense to have AUC as a second objective to be matched, in addition to requiring that the target class 1 prior probability is reached. But even if both of these conditions are met there can still be variation in the mean of the concave function. This is demonstrated by the ‘Platt scaling’ posterior probabilities for which $E_{Q}\bigl{[}\sqrt{\eta_{Q}(X)}\bigr{]}$ takes a notedly lower value than for the other four QMM posterior probabilities with almost identical values.
@@ 6 Conclusions
Recalibration of binary probabilistic classifiers to a target prior probability is an important task in areas like credit risk management. This paper presents analyses and methods for recalibration from a distribution shift perspective. It has turned out that distribution shift assumptions linked to the performance in terms of area under the curve (AUC) of a probabilistic classifier under the source distribution are useful for the design of meaningful recalibration methods. Two new methods called parametric CSPD (covariate shift with posterior drift) and ROC-based QMM (quasi moment matching) have been proposed and have been tested together with some other methods in an example setting. The outcomes of this testing exercise suggest that the QMM methods discussed in the paper can provide appropriately conservative results in evaluations with concave functions like for instance risk weights functions for credit risk.
{credits}
6.0.1 Acknowledgements
The author would like to thank two anonymous reviewers for their useful comments and suggestions.
@@ Appendix 0.A Appendix
0.A.1 The AUC of discrete auto-calibrated probabilitistic classifiers
How to calculate implied AUC for an auto-calibrated [^4] The probabilistic classifier $S$ is auto-calibrated if for all $s$ in the range of $S$ it holds that $\mu[Y=1\,|\,S=s]=s$ (Tsyplakov [26], Section 2.2). probabilistic classifier $S$ under a joint distribution $\mu(S,Y)$ of $S$ and $Y\in\mathcal{Y}=\{0,1\}$? We assume here that the distribution of $S$ under distribution $\mu$ is discrete, i.e. it is is given by pairs of score values $s_{i}$ and their probabilities $\mu[S=s_{i}]=\pi_{i}>0$, $i=1,\ldots,n$. Then it follows from the AUC-definition and properties in Section 3.3 that
@@ 0.A.1 The AUC of discrete auto-calibrated probabilitistic classifiers
0.A.2 Adapting discrete distributions for QMM